How do you figure out the budgeting for a food business chain store. Please discuss briefly with bar and charts if possible otherwise normal report.
Answer #1

As The average food cost percentage is 35% for a fine dining restaurant and 25% for a quick service restaurant. How do you find this number? Look at your weekly inventory.

Food Cost Percentage = (Beginning Inventory + Purchases – Ending Inventory) ÷ Food Sales.

Check out the example below to see this formula in action.

Beginning Inventory= $15,000
Purchases= $4,000
Ending Inventory = $16,000
Food Sales = $10,000 Food Cost Percentage= (15,000 + 4,000 – 16,000) ÷ 10,000
Food Cost Percentage= 3,000 ÷ 10,000
Food Cost Percentage= 0.30, or 30%

  1. List all food supplies you received at the start of week.
    Many inventory management systems work on tablets, so you can walk around the back of house checking off items like you would on a clipboard.

  2. Sum together the dollar value of each item.
    How much did you pay for each box of chicken in your back of house?

  3. Track your purchases.
    Were there any other purchases that you made within the week, after beginning inventory?

  4. Take inventory again at the beginning of next week.
    Follow the same exact process. Many recommend a shelf-to-sheetsystem, where your inventory tracking system is set up as your back-of-house is.

  5. Add together total food sales per shift.
    Many POS systems with restaurant analytics can give you this information automatically.

  6. Calculate actual food cost for the week. Complete the following food cost formula:

Food Cost Percentage = (Beginning Inventory + Purchases – Ending Inventory) ÷ Food Sales.

Check out the example below to see this formula in action.

Beginning Inventory= $15,000
Purchases= $4,000
Ending Inventory = $16,000
Food Sales = $10,000 Food Cost Percentage= (15,000 + 4,000 – 16,000) ÷ 10,000
Food Cost Percentage= 3,000 ÷ 10,000
Food Cost Percentage= 0.30, or 30%